OpenStudy (anonymous):

I need to find all the values of x that are not in the domain of h. f(x)=(x+7)/(x^(2)+14x+49)

jimthompson5910 (jim_thompson5910):

do you mean domain of f ?

OpenStudy (anonymous):

yeah, I do sorry about that.

jimthompson5910 (jim_thompson5910):

the values you want to exclude from the domain are the values that make the denominator zero

jimthompson5910 (jim_thompson5910):

so what values of x make x^2 + 14x + 49 = 0 true?

jimthompson5910 (jim_thompson5910):

Hint: x^2 + 14x + 49 factors to (x+7)(x+7)

OpenStudy (anonymous):

x=7, -7?

jimthompson5910 (jim_thompson5910):

just -7 actually

jimthompson5910 (jim_thompson5910):

(x+7)(x+7) = 0 x+7 = 0 or x+7 = 0 x = -7 or x = -7 x = -7

jimthompson5910 (jim_thompson5910):

so -7 is the only value that makes x^2 + 14x + 49 equal to zero

jimthompson5910 (jim_thompson5910):

that means -7 is the only number that's not in the domain of f

OpenStudy (anonymous):

Okay thanks. It's been years since I've had a math class and this algebra is kicking my rear.

jimthompson5910 (jim_thompson5910):

well just kick it back harder and you'll be fine

OpenStudy (anonymous):

:) Thanks again for the help

jimthompson5910 (jim_thompson5910):

you're welcome